# Lennart Berg and Torben Wiedemann

### v1

@@ 4 Foldings
On the level of Coxeter groups, $(\langle s_{1},s_{2}\rangle,\{s_{1},s_{2}\})$ is a Coxeter system of type $H_{2}$ where $s_{1}\coloneqq\sigma_{e_{1}-e_{2}}\sigma_{e_{3}-e_{4}}$ and $s_{2}\coloneqq\sigma_{e_{2}-e_{3}}\sigma_{e_{4}-e_{5}}$. This realisation of $\operatorname{Weyl}(H_{2})$ as a subgroup of $\operatorname{Weyl}(A_{4})$ is a special case of the construction in [Mü93]: Given a Coxeter system $(W,S)$ and a partition $I=\{I_{1},\ldots,I_{k}\}$ of $S$ which has the so-called property of being admissible, the pair $(\langle R\rangle,R)$ is a Coxeter system where $R=\{r_{1},\ldots,r_{k}\}$ and $r_{j}$ is the longest element in the Coxeter system $(\langle I_{j}\rangle,I_{j})$ for all $j\in\{1,\ldots,k\}$. In all cases we consider, each set $I_{j}$ has cardinality $2$ and its elements $s_{j}$, $t_{j}$ commute (that is, $(\langle I_{j}\rangle,I_{j})$ is of type $A_{1}\times A_{1}$), so that $r_{j}=s_{j}t_{j}$.
We now turn to the realisation of $H_{3}$ as a folding of $D_{6}$. We will find the non-reduced root system $GH_{3}$, the “golden version of $H_{3}$, as an intermediate step between $D_{6}$ and $H_{3}$.
Construction 4.5.
We consider the ordered root base
$\Delta_{D}=(\delta_{1},\ldots,\delta_{6})=(e_{1}-e_{2},e_{2}-e_{3},e_{3}-e_{4},e_{4}-e_{5},e_{5}-e_{6},e_{5}+e_{6})$
of $D_{6}$ and the partition $\{\{\delta_{1},\delta_{6}\},\{\delta_{2},\delta_{4}\},\{\delta_{3},\delta_{5}\}\}$ of $\Delta_{D}$ given by the diagram in Figure 1(b). By [Mü93, 1.1, 5.4], $(\langle s_{1},s_{2},s_{3}\rangle,R)$ is a Coxeter system of type $H_{3}$ where $s_{1}\coloneqq\sigma_{\delta_{1}}\sigma_{\delta_{6}}$, $s_{2}\coloneqq\sigma_{\delta_{2}}\sigma_{\delta_{4}}$, $s_{3}\coloneqq\sigma_{\delta_{3}}\sigma_{\delta_{5}}$ and $R\coloneqq\{s_{1},s_{2},s_{3}\}$.
We briefly sketch how to realise the embedding of $\operatorname{Weyl}(H_{3})$ into $\operatorname{Weyl}(D_{6})$ on the level of root systems. Put $W\coloneqq\langle\delta_{1},\ldots,\delta_{6}\rangle_{\mathbb{R}}$ and define subspaces
@@
$W_{1}$ $\coloneqq\langle\delta_{3}-\tau\delta_{5},\delta_{4}-\tau\delta_{2},\delta_{1}-\tau\delta_{6}\rangle_{\mathbb{R}},\quad W_{2}\coloneqq\langle\delta_{5}+\tau\delta_{3},\delta_{2}+\tau\delta_{4},\delta_{6}+\tau\delta_{1}\rangle_{\mathbb{R}}$
@@
where $\tau\coloneqq\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. A straightforward computation (see 2.11) shows that $W=W_{1}\oplus W_{2}$ is an orthogonal, $R$-invariant decomposition of $W$ and that $\pi(D_{6})$ is a non-reduced root system in $W_{2}$ with $60$ roots where $\pi:W\rightarrow W_{2}$ denotes the orthogonal projection onto $W_{2}$. More precisely, we have $\pi(D_{6})=X\cup\tau X$ where $X$ denotes the set of indivisible roots in $\pi(D_{6})$. The only root systems with 30 roots are $A_{5}$ and $H_{3}$, and an inspection shows that $X$ is of type $H_{3}$. Hence $\pi(D_{6})$ is a root system of type $GH_{3}\coloneqq H_{3}\cup\tau H_{3}$, which we call the golden version of $H_{3}$. Note that the scaling map $\phi\colon v\mapsto\frac{1}{\lVert v\rVert}v$ maps $\pi(D_{6})$ onto $H_{3}$ and that $\pi\colon D_{6}\rightarrow\pi(D_{6})$ is a bijection.
Notation 4.6.
For the rest of this section, we identify $H_{3}$ with the image of $\pi(D_{6})$ under the scaling map $\phi$ in 4.5. Then $\Delta_{H}\coloneqq(\rho_{1},\rho_{2},\rho_{3})$ is a root base of $H_{3}$ in standard order where $\rho_{1}\coloneqq\phi\pi(e_{1}-e_{2})=\phi\pi(e_{5}+e_{6})$, $\rho_{2}\coloneqq\phi\pi(e_{2}-e_{3})=\phi\pi(e_{4}-e_{5})$ and $\rho_{3}\coloneqq\phi\pi(e_{3}-e_{4})=\phi\pi(e_{5}-e_{6})$. We will also denote roots in $H_{3}$ by $[a_{1},a_{2},a_{3}]\coloneqq\sum_{i=1}^{3}a_{i}\rho_{i}$ for $a_{1},a_{2},a_{3}\in\mathbb{R}$. For all $\alpha\in H_{3}$, we denote by $\kappa^{\text{sh}}(\alpha)$ and $\kappa^{\text{lo}}(\alpha)$ the two roots in $(\phi\circ\pi)^{-1}(\alpha)\cap D_{6}$, ordered so that $\pi(\kappa^{\text{lo}}(\alpha))=\tau\cdot\pi(\kappa^{\text{sh}}(\alpha))$ (that is, $\pi(\kappa^{\text{lo}}(\alpha))$ is long in $GH_{3}$ and $\pi(\kappa^{\text{sh}}(\alpha))$ is short). Further, we denote by $\Pi_{D}$, $\Pi_{H}$ and $\Pi_{GH}$ the positive systems in $D_{6}$, $H_{3}$ and $GH_{3}$ corresponding to $\Delta_{D}$, $\Delta_{H}$ and $\pi(\Delta_{D})$, respectively.
$\alpha$ $\kappa^{\text{sh}}(\alpha)$ $\kappa^{\text{lo}}(\alpha)$
@@
$\alpha$ $\kappa^{\text{sh}}(\alpha)$ $\kappa^{\text{lo}}(\alpha)$
@@
Figure 2: The preimages of the positive system in $H_{3}$ under $\phi\circ\pi$ in $D_{6}$. See 4.5 and 4.6.
Observation 4.7.
For all $\alpha\in H_{3}$, the root $\kappa^{\text{lo}}(\alpha)$ is adjacent to $\kappa^{\text{sh}}(\alpha)$ and $-\kappa^{\text{sh}}(\alpha)$ (in $D_{6}$). This follows from an inspection of Figure 2.
Observation 4.8.
We have $\Pi_{D}=\{\kappa^{\text{lo}}(\alpha),\kappa^{\text{sh}}(\alpha)\mid\alpha\in\Pi_{H}\}$, which follows from an inspection of Figure 2. Hence $\Pi_{H}=\phi\pi(\Pi_{D})$.
@@
In a similar way as in 4.5, the orthogonal projection of $A_{4}$ on a suitable subspace yields the non-reduced root system $GH_{2}=H_{2}\cup\tau H_{2}$. See Figure 3 for a graphic depiction of $GH_{2}$ and [MW23] for more details on $GH_{2}$, which inspired our construction of $GH_{3}$.
Now that we have defined maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$, we can give an example of an $H_{3}$-graded group by imitating the construction in 4.3. Along the we, we will also obtain a group with crystallographic $GH_{3}$-commutator relations.
Construction 4.10.
Denote by $e_{ij}\in\mathcal{R}^{6\times 6}$ the matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position. For each root $\alpha\in D_{6}$, we define a root homomorphism $\tilde{\theta}_{\alpha}^{D}\colon(\mathcal{R},+)\rightarrow\operatorname{GL}_{12}(\mathcal{R})$ by the formulas
$\tilde{\theta}_{e_{i}-e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}e_{ij}&0\\ 0&-e_{ji}\end{pmatrix}$
@@
$\tilde{\theta}_{e_{i}+e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}0&e_{ij}-e_{ji}\\ 0&0\end{pmatrix},\quad\tilde{\theta}_{-e_{i}-e_{j}}^{D}(\lambda)\coloneqq\mathord{\mathrm{id}}+\lambda\begin{pmatrix}0&0\\ e_{ji}-e_{ij}&0\end{pmatrix}$

### v2

@@ 4 Foldings
At the level of Coxeter groups, $(\langle s_{1},s_{2}\rangle,\{s_{1},s_{2}\})$ is a Coxeter system of type $H_{2}$ where $s_{1}\coloneqq\sigma_{e_{1}-e_{2}}\sigma_{e_{3}-e_{4}}$ and $s_{2}\coloneqq\sigma_{e_{2}-e_{3}}\sigma_{e_{4}-e_{5}}$. This realisation of $\operatorname{Weyl}(H_{2})$ as a subgroup of $\operatorname{Weyl}(A_{4})$ is a special case of the construction in [Mü93]: Given a Coxeter system $(W,S)$ and a partition $I=\{I_{1},\ldots,I_{k}\}$ of $S$ which has the so-called property of being admissible, the pair $(\langle R\rangle,R)$ is a Coxeter system where $R=\{r_{1},\ldots,r_{k}\}$ and $r_{j}$ is the longest element in the Coxeter system $(\langle I_{j}\rangle,I_{j})$ for all $j\in\{1,\ldots,k\}$. In all cases we consider, each set $I_{j}$ has cardinality $2$ and its elements $s_{j}$, $t_{j}$ commute (that is, $(\langle I_{j}\rangle,I_{j})$ is of type $A_{1}\times A_{1}$), so that $r_{j}=s_{j}t_{j}$.
Remark 4.7.
The same procedure as in 4.6 allows us to realise $\operatorname{Weyl}(H_{3})$ and $\operatorname{Weyl}(H_{4})$ as subgroups of $\operatorname{Weyl}(D_{6})$ and $\operatorname{Weyl}(E_{8})$, respectively. Namely, let $\Delta_{E}=(\delta_{1},\ldots,\delta_{8})$ be an ordered root base of $E_{8}$ as in 4.1. Then by [Mü93, 1.1, 5.4], $(\langle s_{0},s_{1},s_{2},s_{3}\rangle,R)$ is a Coxeter system of type $H_{4}$ where $s_{0}\coloneqq\sigma_{\delta_{7}}\sigma_{\delta_{8}}$, $s_{1}\coloneqq\sigma_{\delta_{1}}\sigma_{\delta_{6}}$, $s_{2}\coloneqq\sigma_{\delta_{2}}\sigma_{\delta_{4}}$, $s_{3}\coloneqq\sigma_{\delta_{3}}\sigma_{\delta_{5}}$ and $R\coloneqq\{s_{0},s_{1},s_{2},s_{3}\}$. Similarly, $(\langle s_{1},s_{2},s_{3}\rangle,\{s_{1},s_{2},s_{3}\})$ is a Coxeter system of type $H_{3}$. These foldings are encoded by the diagrams in Figures 2(b) and 2(c).
Remarks 4.6 and 4.7 describe the folding $X\rightarrow H_{\ell}$ at the level of Coxeter groups – that is, as an embedding $\operatorname{Weyl}(H_{\ell})\rightarrow\operatorname{Weyl}(X)$. To obtain a map $\pi\colon X\rightarrow H_{\ell}$, we will first project onto a suitable subspace and then scale to the unit circle. The first step yields a map $\tilde{\pi}\colon X\rightarrow GH_{\ell}$, which is described in the following proposition.
Proposition 4.8.
The following hold:
  1. (a)
    There exists a unique bijection $\tilde{\pi}\colon E_{8}\rightarrow GH_{4}$ with
    $\tilde{\pi}(\delta_{7})$ $=\rho_{0},$ $\tilde{\pi}(\delta_{1})$ $=\rho_{1},$ $\tilde{\pi}(\delta_{2})$ $=\rho_{2},$ $\tilde{\pi}(\delta_{3})$ $=\tau\rho_{3},$
    $\tilde{\pi}(\delta_{8})$ $=\tau\rho_{0},$ $\tilde{\pi}(\delta_{6})$ $=\tau\rho_{1},$ $\tilde{\pi}(\delta_{4})$ $=\tau\rho_{2},$ $\tilde{\pi}(\delta_{5})$ $=\rho_{3}$
    and such that $\tilde{\pi}$ is induced by a linear map $\langle E_{8}\rangle_{\mathbb{R}}\rightarrow\langle GH_{4}\rangle_{\mathbb{R}}$ on the surrounding vector spaces.
  2. (b)
    The map $\tilde{\pi}$ satisfies $\tilde{\pi}(D_{6})=GH_{3}$ and $\tilde{\pi}(A_{4})=GH_{2}$. Here $D_{6}$, $A_{4}$ and $GH_{3}$, $GH_{2}$ are embedded into $E_{8}$, $GH_{4}$, respectively, as in 4.1.
  3. (c)
    We have
    $\tilde{\pi}(\alpha^{\sigma(\delta_{i})\sigma(\delta_{j})})=\tilde{\pi}(\alpha)^{\sigma(\tilde{\pi}(\delta_{i}))}=\tilde{\pi}(\alpha)^{\sigma(\tilde{\pi}(\delta_{j}))}$
    for all $\alpha\in E_{8}$ and all $(i,j)\in\{(7,8),(1,6),(2,4),(5,3)\}$.
  4. (d)
    For each $w\in\operatorname{Weyl}(H_{4})$, there exists a unique $u(w)\in\operatorname{Weyl}(E_{8})$ with $\tilde{\pi}(\alpha^{u(w)})=\tilde{\pi}(\alpha)^{w}$ for all $\alpha\in E_{8}$. This defines an injective homomorphism $\operatorname{Weyl}(H_{4})\rightarrow\operatorname{Weyl}(E_{8}),w\mapsto u(w)$ which maps $\operatorname{Weyl}(H_{3})$ into $\operatorname{Weyl}(D_{6})$.
  5. (e)
    The map $\pi\coloneqq\phi\circ\tilde{\pi}\colon E_{8}\rightarrow H_{4}$ (where $\phi\colon GH_{4}\rightarrow H_{4}$ denotes the scaling map from 2.16) is a surjection with $\lvert\pi^{-1}(\beta)\rvert=2$ for all $\beta\in H_{4}$ and $\tilde{\pi}(\alpha^{u(w)})=\tilde{\pi}(\alpha)^{w}$ for all $\alpha\in E_{8}$ and $w\in\operatorname{Weyl}(H_{4})$.
Proof.
@@
Put $W\coloneqq\langle\delta_{1},\ldots,\delta_{8}\rangle_{\mathbb{R}}$ and define subspaces
$W_{1}$ $\coloneqq\langle\delta_{8}-\tau\delta_{7},\delta_{6}-\tau\delta_{1},\delta_{4}-\tau\delta_{2},\delta_{3}-\tau\delta_{5}\rangle_{\mathbb{R}},$
$W_{2}$ $\coloneqq\langle\delta_{7}+\tau\delta_{8},\delta_{1}+\tau\delta_{6},\delta_{2}+\tau\delta_{4},\delta_{5}+\tau\delta_{3}\rangle_{\mathbb{R}}$
@@
where $\tau\coloneqq\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. Denote by $\tilde{\pi}:W\rightarrow W_{2}$ the orthogonal projection onto $W_{2}$ and define $R$ as in 4.7. A straightforward computation (see 2.18) shows that $W=W_{1}\oplus W_{2}$ is an orthogonal, $R$-invariant decomposition of $W$, that $\tilde{\pi}$ is injective on $E_{8}$ (so that $\lvert\tilde{\pi}(E_{8})\rvert=240$) and that $\tilde{\pi}(E_{8})$ is a non-reduced root system in $W_{2}$. More precisely, we have $\tilde{\pi}(E_{8})=Y\cup\tau Y$ where $Y$ denotes the set of indivisible roots in $\tilde{\pi}(E_{8})$. We conclude that $Y$ is of type $H_{4}$ because this is the only reduced root system of rank at most $4$ with $120$ roots. Hence $\tilde{\pi}(E_{8})$ is (up to scaling) a root system of type $GH_{4}$. Another computation shows that $(\tilde{\pi}(\delta_{7}),\tilde{\pi}(\delta_{1}),\tilde{\pi}(\delta_{2}),\tilde{\pi}(\delta_{5}))$ is a root base of $\tilde{\pi}(E_{8})$ (in the sense of 2.16) and that $\tilde{\pi}(\delta_{j})=\tau\tilde{\pi}(\delta_{i})$ for all $(i,j)\in\{(7,8),(1,6),(2,4),(5,3)\}$. By identifying $\tilde{\pi}(E_{8})$ with $GH_{4}$ via the root bases $(\tilde{\pi}(\delta_{7}),\tilde{\pi}(\delta_{1}),\tilde{\pi}(\delta_{2}),\tilde{\pi}(\delta_{5}))$ and $(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$, we obtain a map $E_{8}\rightarrow GH_{4}$ which is induced by a linear map and which acts on $(\delta_{1},\ldots,\delta_{8})$ in the desired way. Since a linear map is determined by its action on a basis, $\tilde{\pi}$ is necessarily unique. This proves a.
Since $\tilde{\pi}(\{\delta_{1},\ldots,\delta_{6}\})\subseteq GH_{3}$ and $\tilde{\pi}$ is (induced by a) linear map, we have $\tilde{\pi}(D_{6})\subseteq GH_{3}$. Now let $\beta\in GH_{3}$. Since $\tilde{\pi}$ is bijective, there exists $\alpha\in E_{8}$ with $\tilde{\pi}(\alpha)=\beta$. Since $(\delta_{1},\ldots,\delta_{8})$ is a root base and $\tilde{\pi}$ is linear, we may without loss of generality assume that $\alpha$ is positive and write $\alpha=\sum_{i=1}^{8}a_{i}\delta_{i}$ with $a_{1},\ldots,a_{8}\in\mathbb{Z}_{>0}$. Then
$\beta=\tilde{\pi}(\alpha)$ $=(a_{7}+\tau a_{8})\rho_{0}+(a_{1}+\tau a_{6})\rho_{1}+(a_{2}+\tau a_{4})\rho_{2}+(a_{5}+\tau a_{3})\rho_{3}.$
Since $\beta\in GH_{3}$, we conclude that $a_{7}=-\tau a_{8}$. This implies that $a_{7}=a_{8}=0$. Thus $\alpha\in D_{6}$, so $\tilde{\pi}(D_{6})=GH_{3}$. In a similar way, one can show $\tilde{\pi}(A_{4})=GH_{2}$. This proves b.
For c, let $(i,j)\in\{(7,8),(1,6),(2,4),(5,3)\}$, let $\alpha\in E_{8}$ and put $r\coloneqq\sigma(\delta_{i})\sigma(\delta_{j})$. Note that the map $\tilde{\pi}$ in our explicit construction is an orthogonal projection, so that $\tilde{\pi}(\alpha^{r})=\tilde{\pi}(\alpha)^{r}$. It remains to check that $r$, $\sigma(\tilde{\pi}(\delta_{i}))$ and $\sigma(\tilde{\pi}(\delta_{j}))$ act identically on any root in $GH_{4}$, which is a straightforward computation (see, again, 2.18).
For d, let $w\in H_{4}$. The existence of $u(w)$ follows from c because $\operatorname{Weyl}(H_{4})$ is generated by $\{\sigma(\rho_{i})\mid i\in\{0,1,2,3\}\}=\{\sigma(\tilde{\pi}(\delta_{j}))\mid j\in\{1,\ldots,8\}\}$. Since $\tilde{\pi}$ is bijective, $u(w)$ is uniquely determined by the desired property and $w\mapsto u(w)$ is an injective homomorphism. Further, our construction shows that $u(w)\in\operatorname{Weyl}(D_{6})$ if $w\in\operatorname{Weyl}(H_{3})$.
Finally, assertion e follows from the fact that $\phi$ is surjective with fibers of cardinality 2 and that $\phi(\alpha^{w})=\phi(\alpha)^{w}$ for all $\alpha\in GH_{4}$ and $w\in\operatorname{Weyl}(H_{4})$. ∎
Note 4.9.
The ordering of $\delta_{1},\ldots,\delta_{8}$ in 4.8 a corresponds to the way the diagram in Figure 2(c) is drawn. Further, the monomorphism $w\mapsto u(w)$ in 4.8 d is precisely the embedding of Weyl groups in 4.7. Note also that the choice of $\tilde{\pi}$ depends on the fixed choices of the root bases $(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ and $(\delta_{1},\ldots,\delta_{8})$.
Note 4.10.
The generators of $W_{1}$ in the proof of 4.8 encode precisely the relations $\delta_{8}\equiv\tau\delta_{7}$, $\delta_{6}\equiv\tau\delta_{1}$, $\delta_{4}\equiv\tau\delta_{2}$ and $\delta_{3}\equiv\tau\delta_{5}$ that we want to realise modulo $\tilde{\pi}$. Further, note that the orbit of $(\delta_{1},\delta_{6})$ under $u(\operatorname{Weyl}(H_{4}))$ contains $(\delta_{2},\delta_{4})$, $(\delta_{5},\delta_{3})$ and $(\delta_{7},\delta_{8})$, as can be seen from the following diagram:
${{(\delta_{7},\delta_{8})}}$${{(\delta_{1},\delta_{6})}}$${{(\delta_{2},\delta_{4})}}$${{(\delta_{5},\delta_{3})}.}$$\scriptstyle{r_{1}r_{0}}$$\scriptstyle{r_{2}r_{1}}$$\scriptstyle{r_{2}r_{3}r_{2}r_{3}}$
Together with the observation from 4.8 that $W_{1}$ is $R$-invariant, it follows that $W_{1}$ is actually the smallest $R$-invariant subspace of $W$ containing any of its four generators. In particular, it is not possible to choose generators of $W_{1}$ so that $\tilde{\pi}(\delta_{2})$ and $\tilde{\pi}(\delta_{3})$ have the same length in $GH_{4}$.
Remark 4.11.
We can also obtain bijections $D_{6}\rightarrow GH_{3}$ and $A_{4}\rightarrow GH_{2}$ as well as surjections $D_{6}\rightarrow H_{3}$ and $A_{4}\rightarrow H_{2}$ with similar properties by restricting $\tilde{\pi}\colon E_{8}\rightarrow GH_{4}$ and $\pi\colon E_{8}\rightarrow H_{4}$ to the subsystems $\Psi_{\delta_{1},\ldots,\delta_{6}}$ and $\Psi_{\delta_{2},\ldots,\delta_{5}}$ of $E_{8}$.
Remark 4.12.
Figures 10 and 11 in the appendix depict all tuples $(\beta,\alpha_{1},\alpha_{2})\in H_{4}\times E_{8}\times E_{8}$ with $\beta$ positive, $\phi\tilde{\pi}(\alpha_{1})=\beta=\phi\tilde{\pi}(\alpha_{2})$ and $\tilde{\pi}(\alpha_{2})=\tau\tilde{\pi}(\alpha_{1})$. (Note that then $(-\beta,-\alpha_{1},-\alpha_{2})$ is a tuple with the same properties except that $\beta$ is negative.) Here we use Notation 2.10 to depict roots via their coordinates with respect to the root bases $(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ and $(\delta_{1},\ldots,\delta_{8})$. For the reader’s convenience, the subtable consisting of all tuples $(\beta,\alpha_{1},\alpha_{2})\in H_{3}\times D_{6}\times D_{6}$ is given in Figure 3.
$\beta$ $\alpha_{1}$ $\alpha_{2}$
@@
$\beta$ $\alpha_{1}$ $\alpha_{2}$
@@
Figure 3: The preimages of the positive roots in $H_{3}$ under $\pi$. See 4.12.
Notation 4.13.
For the rest of this section and in addition to 4.1, we fix the bijection $\tilde{\pi}\colon E_{8}\rightarrow GH_{4}$ and the surjection $\pi\colon E_{8}\rightarrow H_{4}$ from 4.8. If $X\in\{A_{4},D_{6}\}$, we will always consider the restrictions of $\pi$ and $\tilde{\pi}$ to $X$ in place of $\pi$ and $\tilde{\pi}$, as described in 4.11.
Before we can use 4.8 to construct $H_{\ell}$-graded groups, we collect some more useful properties of $\pi$.
Lemma 4.14.
The Weyl group of $H_{\ell}$, embedded into $\operatorname{Weyl}(X)$ via 4.8 d, acts transitively on the set $\{\pi^{-1}(\beta)\mid\beta\in H_{\ell}\}$. More precisely, $\pi^{-1}(\beta^{w})=\pi^{-1}(\beta)^{u(w)}$ for all $w\in\operatorname{Weyl}(H_{\ell})$.
Proof.
Let $\beta\in H_{\ell}$, $w\in\operatorname{Weyl}(H_{\ell})$ be arbitrary. Then $\pi^{-1}(\beta)^{u(w)}\subseteq\pi^{-1}(\beta^{w})$ by 4.8 e. As both sets have cardinality $2$, we conclude that $\pi^{-1}(\beta)^{u(w)}=\pi^{-1}(\beta^{w})$. It follows that $\operatorname{Weyl}(H_{\ell})$ acts on $\{\pi^{-1}(\beta)\mid\beta\in H_{\ell}\}$, and this action is transitive by 2.14. ∎
Lemma 4.15.
For all $\beta\in H_{\ell}$, the preimage $\pi^{-1}(\beta)$ is of the form $\{\alpha_{1},\alpha_{2}\}$ for unique orthogonal roots $\alpha_{1},\alpha_{2}\in X$ with $\tilde{\pi}(\alpha_{2})=\tau\tilde{\pi}(\alpha_{1})$.
Proof.
Let $\beta\in H_{\ell}$. Since $\tilde{\pi}$ is bijective and $\pi=\phi\circ\tilde{\pi}$, we have $\pi^{-1}(\beta)=\{\alpha_{1},\alpha_{2}\}$ for unique distinct roots $\alpha_{1},\alpha_{2}\in X$ with $\tilde{\pi}(\alpha_{2})=\tau\tilde{\pi}(\alpha_{1})$. If $\beta\in\{\rho_{1},\rho_{2},\rho_{3}\}$ (or $\beta=\rho_{0}$ if $\ell=4$), then $\alpha_{1}$, $\alpha_{2}$ are given by  4.8 a, and we can observe that they are orthogonal. Now it follows from 4.14 that $\alpha_{1},\alpha_{2}$ are orthogonal for arbitrary $\beta\in H_{\ell}$. ∎
@@
In 4.15, we implicitly introduce an order on $\pi^{-1}(\beta)$ by declaring that $\alpha_{1}$ is the root in $\pi^{-1}(\beta)$ whose image in $GH_{\ell}$ is short. It is compatible with the action of $u(\operatorname{Weyl}(H_{4}))$: If $\pi(\alpha)$ is short in $GH_{3}$, then so is $\pi(\alpha^{u(w)})=\pi(\alpha)^{w}$ (for $\alpha\in E_{8}$ and $w\in\operatorname{Weyl}(H_{4})$).
We will use this order in 4.21 to decide which factor of a root group will be the “left one” and which one the “right one”. For the computations in 4.27 (specifically, in (4.28)), it will be crucial that this choice is performed in a way that is compatible with the action of the Weyl elements.
Lemma 4.17.
Let $\beta\in H_{4}$ and let $\alpha_{1},\alpha_{2}$ denote the two roots in $\pi^{-1}(\beta)$. Then $u(\sigma_{\beta})=\sigma_{\alpha_{1}}\sigma_{\alpha_{2}}=\sigma_{\alpha_{2}}\sigma_{\alpha_{1}}$.
Proof.
If $\beta\in\{\rho_{0},\ldots,\rho_{3}\}$, the assertion holds by 4.8 c. For arbitrary $\beta$, there exists $w\in\operatorname{Weyl}(H_{4})$ with $\beta=\rho_{1}^{w}$ by 2.14. Hence, by 2.5, $\sigma_{\beta}=\sigma_{\rho_{1}}^{w}$ and
$u(\sigma_{\beta})=u(\sigma(\rho_{1}))^{u(w)}=(\sigma_{\delta_{1}}\sigma_{\delta_{6}})^{u(w)}=\sigma_{\delta_{1}^{u(w)}}\sigma_{\delta_{6}^{u(w)}}.$
Since
$\{\alpha_{1},\alpha_{2}\}=\pi^{-1}(\beta)=\pi^{-1}(\rho_{1}^{w})=\pi^{-1}(\rho_{1})^{u(w)}=\{\delta_{1},\delta_{6}\}^{u(w)}$
by 4.14, we infer that $u(\sigma_{\beta})=\sigma_{\alpha_{1}}\sigma_{\alpha_{2}}$. ∎
We can now turn to the folding procedure for root graded groups.
Notation 4.18.
For the rest of this section, we denote by $G$ an arbitrary group and by $(U_{\alpha}^{X})_{\alpha\in X}$ an $X$-grading of $G$.
Definition 4.19.
The folding of $(U_{\alpha}^{X})_{\alpha\in X}$ (with respect to the partition given by Figure 2) is the family $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ defined by
$U_{\beta}^{H}\coloneqq U_{\pi^{-1}(\{\beta\})}^{X}=\langle U_{\alpha}^{X}\mid\alpha\in X,\pi(\alpha)=\beta\rangle$
for all $\beta\in H_{\ell}$.
In 4.32, we will show that the folding $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ is an $H_{\ell}$-grading of $G$.
Remark 4.20.
For any root $\beta\in H_{\ell}$, the two roots $\alpha_{1},\alpha_{2}$ in $\pi^{-1}(\beta)$ are orthogonal by 4.15 and hence adjacent. Hence $U_{\beta}^{H}=U_{\alpha_{1}}^{X}U_{\alpha_{2}}^{X}=U_{\alpha_{2}}^{X}U_{\alpha_{1}}^{X}$, and this group is abelian. In particular, the construction in 4.3 is a special case of 4.19.
In the examples of Chevalley groups, each root group is parametrised by the additive group of a commutative ring $\mathcal{R}$. This leads to a natural parametrisation of the root groups in the folding by $\mathcal{R}\times\mathcal{R}$.
Definition 4.21.
Let $(\mathcal{R},+)$ be a group and let $(\theta_{\alpha}^{X}\colon\mathcal{R}\rightarrow U_{\alpha}^{X})_{\alpha\in X}$ be a family of root isomorphisms (in the sense of 3.5). The folding of $(\theta_{\alpha}^{X}\colon\mathcal{R}\rightarrow U_{\alpha}^{X})_{\alpha\in X}$ (with respect to the partition given by Figure 2) is the family $(\theta_{\beta}^{H}\colon\mathcal{R}\times\mathcal{R}\rightarrow U_{\beta}^{H})_{\beta\in H_{\ell}}$ defined by
$\theta_{\beta}^{H}(a,b)\coloneqq\theta_{\alpha_{1}}^{X}(a)\theta_{\alpha_{2}}^{X}(b)$
for all $\beta\in H_{\ell}$ where $(\alpha_{1},\alpha_{2})$ is the unique pair of roots in $X$ with $\pi(\alpha_{1})=\beta=\pi(\alpha_{2})$ and $\tilde{\pi}(\alpha_{2})=\tau\tilde{\pi}(\alpha_{1})$.
Note that the group $(\mathcal{R},+)$ in 4.21 is necessarily abelian by 3.6.
In the following, we give two more examples of $X$-graded groups, which give rise to foldings in the sense of 4.19 and 4.21. They are equipped with families $(\theta_{\alpha}^{X,\text{n}})_{\alpha\in X}$ of root isomorphisms where the letter “n” stands for “naive”. In 4.26, we will define twists of these families which ensure that certain structure constants (or rather, signs) behave in a specific way.
Example 4.22.
Let $\mathcal{R}$ be a commutative ring and denote by $e_{ij}\in\mathcal{R}^{6\times 6}$ the matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position. For each root $\alpha\in D_{6}$, we define a root homomorphism $\tilde{\theta}_{\alpha}^{D,\text{n}}\colon(\mathcal{R},+)\rightarrow\operatorname{GL}_{12}(\mathcal{R})$ by the formulas
$\tilde{\theta}_{e_{i}-e_{j}}^{D,\text{n}}(\lambda)\coloneqq\operatorname{id}+\lambda\begin{pmatrix}e_{ij}&0\\ 0&-e_{ji}\end{pmatrix}$
@@
$\tilde{\theta}_{e_{i}+e_{j}}^{D,\text{n}}(\lambda)\coloneqq\operatorname{id}+\lambda\begin{pmatrix}0&e_{ij}-e_{ji}\\ 0&0\end{pmatrix},\quad\tilde{\theta}_{-e_{i}-e_{j}}^{D,\text{n}}(\lambda)\coloneqq\operatorname{id}+\lambda\begin{pmatrix}0&0\\ e_{ji}-e_{ij}&0\end{pmatrix}$